Grade 11Wave Optics

A nearsighted person cannot clearly see beyond 200 cm. Find the power of the lens needed to see objects at large distance.

Radhika Batra
12 Years agoGrade 11

1 Answer

Kevin Nash
12 Years ago
Sol. For the near sighted person, u = ∞ and v = – 200 cm = – 2m So, 1/f = 1/v – 1/u = 1/-2 – 1/∞ = - 0.5 So, power of the lens is –0.5D